OpenStudy (lovelyanna):

Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form. A. (y – 1) = 3(x + 2) B. (x – 2) = 3(y + 1) C. (x + 2) = 3(y – 1) D. (y + 1) = 3(x – 2)

OpenStudy (lovelyanna):

C?

OpenStudy (lovelyanna):

@jigglypuff314 Is C correct?

OpenStudy (lovelyanna):

Ok, thank you

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)

OpenStudy (lovelyanna):

So, C is not correct?

jigglypuff314 (jigglypuff314):

nope

OpenStudy (lovelyanna):

Hmm.. than B?

OpenStudy (lovelyanna):

then*

jigglypuff314 (jigglypuff314):

it would start with a y

OpenStudy (lovelyanna):

Oh.. I still don't get it.. all we have left is A and D, and I might say A, but what if D is correct. Can you explain more?

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) so given slope=m=3 and point (2, -1) plug in 2 = x1 and -1=y1

OpenStudy (lovelyanna):

D?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Phew :) thanks!

jigglypuff314 (jigglypuff314):

you're welcome ;)

OpenStudy (lovelyanna):

:)